7k+ar=-6y+r solve for r

Answer by RAY100(1637)

7k +ar = -6y +r,,,,,subt (ar) and adding(6y) to both sides
.
7k + 6y = r - ar,,,,,,factor
.
7k+6y = r(1-a),,,,,,divide by (1-a)
.
(7k+6y) / (1-a) = r
.